a uniform sphere has a radius of 0.10 m. the moment of inertia of the sphere when axis


of rotation is located at its center is 0.08 kg-m2


, find the mass of the sphere.



1
Expert's answer
2022-04-17T17:08:30-0400
"J=\frac{2}{5}mR^2"

"m=\frac{5J}{2R^2}=\frac{5*0.08}{2*0.10^2}=20\:\rm kg"